# Optimal Compressed Sensing for Image Reconstruction with Diffusion Probabilistic Models

### v2

@@ 3.3 Optimized measurements for CelebA
To test the generality of our method for different image classes, we repeated our experiments on the CelebA dataset (Liu et al., 2015), which consists of about $200,000$ centered face images. We resized all images to $80\times 80$ and converted them to grayscale. The rest of the procedure is the same as described above. The bottom half of Figure 2 shows an example face image from the test set. Similar to the MNIST case, we observe that using the denoiser prior to reconstruct from PCs significantly improves the linear inverse estimates. Notably, the conditional samples all appear to be realistic face images even for low $k$; this indicates that the denoiser prior adequately captured the statistical structure of image datasets. As $k$ increases, the reconstructed face images increasingly resemble the originals. Optimizing the measurement matrix offers further improvements in the results, in this case most visible for $k=16$. See Supplemental Figure 3 for two more examples, including reconstructions obtained with random measurements.
The effect of the optimized measurements is investigated quantitatively in Figure 3B. As with the MNIST dataset, we observe that OLM leads to superior reconstruction using the denoiser prior, illustrated by the red curve in the left panel. Note, however, that the improvements in terms of PSNR is smaller than for the MNIST case, although the improvement is noticeable and consistent across all values of $k$. The peak difference in this case is 0.44 dB in PSNR at $k=2048$. Thus, the possible performance gain between PCA and OLM also depends on the specific image dataset. Investigating what aspects of the image statistics drive this difference is of great interest for future study. See also Supplemental Table 1 for comparison of our results to other methods.
Consistent with results on MNIST, for PCs, reconstruction using denoiser prior results in higher performance than linear reconstruction, and all of these methods outperform reconstruction from random measurements. We confirmed that the improvements from the OLM are nearly universal for all images in our test set (Figure 3B, right panel). For completeness, we also used independent component analysis (ICA, Hyvärinen & Oja 2000) to generate linear measurement vectors. On both datasets, we found the ICA measurement performed marginally worse compared to the PCs.
@@ 3.4 Characterizing the optimized linear measurements
Figure 4: Optimized measurement vectors are distinct from PC and Random. A) Left: Twenty-five selected principal components of the MNIST dataset are visualized as images. They are ordered by variance explained from top to bottom, and left to right. Right: Twenty-five components from the columns of the OLM for $k=64$. B) Same as A) but for the CelebA dataset. C) The variance of the measurement on each component as a function of number of measurements ($k$) for the PCs (blue) and OLMs (red). Note that the optimal measurements are sorted post hoc based on the variance explained by each component. These vectors are optimized jointly for each $k$ and are not ordered by the algorithm itself. D) The Grassmann distance between subspaces spanned by different linear measurements as a function of $k$ (Hamm & Lee, 2008). We show the distance between subspaces spanned by PCs obtained on two random halves of the training data (blue), between subspaces spanned by PCs and OLM (blue), and between subspaces spanned by PCs and a set of random measurements (green). See Appendix E for the definition of Grassmann distance. E) The skewness of the measurement on each component as a function of number of measurements ($k$, sorted post hoc based on the skewness value) for the PCs (blue) and OLMs (red). F) Distribution of mutual information between all pairs of measurement directions for PCA (blue), ICA (purple), and OLM (red).
@@
In this section, we present some qualitative and quantitative analyses to help interpret the difference between the linear measurement subspaces defined by PCs and OLMs. For the MNIST dataset, as expected, the first few PCs appear digit-like, and the measurement vectors contain increasing high-spatial frequency content as $k$ increases (Figure 4A, left). The OLMs, on the other hand, do not follow a coarse to fine or a low to fine frequency ordering and appear to have similar patterns (Figure 4A, right). This is reflected, quantitatively, in the exponential decay in measurement variance as a function $k$ for the PCs, while for the OLMs the measurement variance is relatively constant (Figure 4C, left).
@@
Similar phenomena are observed for the CelebA dataset (Figure 4B, C). As expected, the PCs in this case are “eigenfaces” (Turk & Pentland, 1991): they contain features that are geometrically aligned with features of the face, with increasing spatial frequency content. The variance of the measurements across PCs fall exponentially. Although the OLMs also have a face-like appearance, they differ from the PCs. Each OLM vector resembles a distorted face with a different identity, giving them an overall template-like appearance. Like the MNIST case, the OLMs all appear to have similar spatial frequency compositions. This is consistent with the variances of the measurements, which fall quite slowly across the OLMs (Figure 4C, right).
@@
We can also characterize the linear measurements by analyzing the distribution of measurement values they produce in response to natural images (i.e., the measurement distribution). We found that OLMs differ from traditional methods in two key aspects: First, they generate measurement distributions that are highly asymmetrical, as indicated by the significant non-zero skewness values (see Figure 4E). Second, unlike PCA or ICA, OLMs produce non-factorized measurement distributions, where the measurement axes are not fully independent. This dependency is reflected in the increased pairwise mutual information between different measurement directions observed for OLMs compared to those obtained with PCA or ICA (Figure 4F). This observation aligns with the intuition from the 2D example in Figure 1, where optimal measurements for non-Gaussian priors are not necessarily aligned with the axes.
@@ 3.5 Optimized linear measurements for perceptual loss
Figure 5: Optimal measurement for perceptual loss (structural similarity index measure - SSIM) (Wang et al., 2004). A) Left: SSIM as a function of number of measurement $k$, for PC projection, PC, and OLM optimized for MSE and SSIM, all three using the denoiser prior. The number indicates the increase in SSIM from OLM optimized for MSE to OLM optimized for SSIM. Right: Scatter plot of the SSIM value for individual images, for OLM optimized for SSIM (x-axis) and MSE (y-axis), respectively. B) Example linear inverse solutions obtained from different measurement matrices combined with the denoiser prior. The numbers indicate the SSIM value for each individual image. C) Example measurement vectors from the OLM optimized for SSIM.
The previous sections described measurements optimized for MSE. However, our method is general, and can be used to optimize any differentiable objective function defined on the linear inverse estimates (with the obvious caveat that we cannot guarantee optimality if the function is non-convex).
As an example, we compute a set of OLMs with respect to structural similarity index (SSIM) (Wang et al., 2004), a commonly-used perceptual image quality metric. Figure 5 shows results obtained using this set of optimal linear measurements. We obtained a noticeable improvement in mean SSIM for a range of $k$, and this improvement applies to the majority of the images in the test set. In Figure 5B, four example images are shown to highlight visually the improvement obtained by optimizing for the SSIM objective. Lastly, in Figure 5C, the SSIM optimized measurement vectors are visualized. These are distinct from both the PCs and MSE optimized vectors shown in Figure 4.
@@ 4 Related Work
Image Priors. Traditionally, prior models have been developed by combining constraints imposed by structural properties, such as translation or dilation invariance, with simple parametric forms, such as Gaussian, mixtures of Gaussian, or local Markov random fields. These models have been used for solving inverse problems with steady and gradual improvement in performance (Donoho, 1995; Simoncelli & Adelson, 1996; Moulin & Liu, 1999; Hyvärinen, 1999; Romberg et al., 2001; Şendur & Selesnick, 2002; Portilla et al., 2003a; Lyu & Simoncelli, 2008). More recently, methods such as VAEs (Kingma & Welling, 2013) and GANs (Goodfellow et al., 2020) have been developed to learn more complex image priors by taking advantage of the increased expressivity of deep neural networks. In the last few years, score-based diffusion models have emerged as the state-of-the-art method for learning sophisticated image priors (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Nichol & Dhariwal, 2021), as evidenced by the high quality of generated images obtained as draws from these priors. Diffusion models rely on the explicit relationship between the MMSE denoising solution and the score of the noisy image density (Miyasawa, 1961; Raphan & Simoncelli, 2011; Vincent, 2011). In addition to enabling unconditional sampling from the implicit prior, the power of that prior can be utilized to obtain high-quality solutions to inverse problems (Kadkhodaie & Simoncelli, 2020; Kawar et al., 2022; Chung et al., 2022). A closely related line of work known as plug-and-play (P&P) (Venkatakrishnan et al., 2013) used denoisers as regularizers for solving inverse problems. A number of recent extensions have used this concept to develop MAP solutions for inverse problems (Chan et al., 2016; Romano et al., 2017a; Zhang et al., 2017; Kamilov et al., 2017; Meinhardt et al., 2017; Chan et al., 2017; Mataev et al., 2019; Teodoro et al., 2019; Sun et al., 2019; Reehorst & Schniter, 2019; Pang et al., 2020; Sun et al., 2023).
@@
Optimized Linear Measurements. Other work has considered the problem of optimizing linear measurements directly. Weiss et al. (2007) observed that neither PCA nor random measurement fully takes into account the statistics of natural images. They attempted to optimize linear measurement using information-based criteria, but were not able to find measurements that outperformed PCA. Wu et al. (2019) also proposed a method to find better measurement functions (i.e., both linear and nonlinear) in a compressed sensing framework by using the restricted isometry property (Candes, 2008) as an objective. They showed that the optimized measurement functions (both linear and nonlinear) are superior to simple random measurements, but did not compare their results with PCA. In a different line of work, Burge & Geisler (2015) developed a method for finding linear measurements that are optimal for specific downstream tasks, such as estimating the motion presented in a video sequence.
@@ 5 Conclusion
We present a method for finding optimal linear measurements, using a nonlinear reconstruction method based on a learned prior embedded in a denoiser. We show numerically that the set of linear measurements found through our method, result in superior image reconstruction. This result shows that for signals with non-Gaussian distributions, better measurements exist for minimizing MSE, even though projections onto the PCs maximize the explained variance. The key components of our method are: (i) a denoising diffusion model, which allows us to learn a complex prior underlying the datasets; (ii) a constrained sampling algorithm for obtaining linear inverse estimates from the diffusion prior; and (iii) an end-to-end optimization procedure to find the measurement matrix that minimizes a loss function defined with respect to the linear inverse estimates. We show that the optimal measurement are sensitive to both the statistics of the image dataset, and to the objective function for which they are optimized. Our results highlight the importance of accurately modeling the statistics of the signals to design efficient linear measurements.
One major limitation of our current work is that optimizing linear measurements through the iterative diffusion process is more computationally expensive than PCA and CS. Additionally, the OLMs are obtained separately for each $k$, and are therefore not sequentially ordered like the principal components. Here, we have restricted ourselves to linear measurements to maintain a direct connection to classical signal processing literature, as they are easier to understand and visualize. An interesting direction for future work is to analyze in greater depth the mechanisms through which the measurement subspace specified by OLMs leads to better linear inverse estimates. Additionally, it is of interest to expand the current results to linear measurements with realistic noise models and measurement constraints such as locality and non-negativity. Finally, extending our work to nonlinear measurement functions will be of great interest, representing a form of nonlinear compression problem (e.g., Ballé et al. (2016)), with reconstruction through conditional sampling from diffusion models.
@@ Appendix B Principal Components Analysis
To minimize $l(M)$, the $m_{i}$’s should be the eigenvectors associated with the largest eigenvalues of $XX^{T}$, which is the empirical covariance matrix. Note that this solution does not assume a Gaussian data distribution: For any data, the optimal linear projection is the projection onto the leading principal components of the covariance matrix.
When the data is in fact Gaussian distributed, linear projection onto the first $k$ PCs does achieve the minimal error possible, as shown in the bivariate example in Fig. 1A. Otherwise, for non-Gaussian distribution, nonlinear estimation based on the linear measurements can be used to improve performance.

### v3

@@ 3.3 Optimized measurements for CelebA
To test the generality of our method for different image classes, we repeated our experiments on the CelebA dataset (Liu et al., 2015), which consists of about $200,000$ centered face images. We resized all images to $80\times 80$ and converted them to grayscale. The rest of the procedure is the same as described above. The bottom half of Figure 2 shows an example face image from the test set. Similar to the MNIST case, we observe that using the denoiser prior to reconstruct from PCs significantly improves the linear inverse estimates. Notably, the conditional samples all appear to be realistic face images even for low $k$; this indicates that the denoiser prior adequately captured the statistical structure of this image dataset. As $k$ increases, the reconstructed face images increasingly resembles the original. Optimizing the measurement matrix offers further improvements in the results, in this case most visible for $k=16$. See Supplemental Figure 4 for two more examples, including reconstructions obtained with random measurements.
The effect of the optimized measurements is quantified in Figure 3B. As with the MNIST dataset, we observe that OLM leads to superior reconstruction using the denoiser prior, illustrated by the red curve in the left panel. The improvements arising from OLM measurements are consistent across all values of $k$, although they are smaller than for the MNIST case. The peak difference in this case is 0.44 dB in PSNR at $k=2048$. Thus, the performance gain between PCA and OLM also depends on the specific image dataset.
Consistent with results on MNIST, reconstruction from PCs using the denoiser prior results in higher performance than linear reconstruction, and all of these methods outperform reconstruction from random measurements. We confirmed that the improvements from the OLM are nearly universal for all images in our test set (Figure 3B, right panel). For completeness, we also used independent component analysis (ICA, Hyvärinen & Oja 2000) to generate linear measurement vectors. On both datasets, we found that reconstructions using the denoiser prior based on ICA measurements were marginally worse than those using the PCs (not shown).
@@ 3.4 Characterizing the optimized linear measurements
Figure 4: Optimized measurement vectors are distinct from PC and random measurements. A) Left: Twenty-five selected principal components of the MNIST dataset, visualized as images, sorted by variance explained from top to bottom, and left to right. Right: Twenty-five components from the columns of the OLM for $k=64$. B) Same as A) but for the CelebA dataset. C) The variance of the measurement on each component as a function of number of measurements ($k$) for the PCs (blue) and OLMs (red). Note that the optimal measurements are sorted post hoc based on the variance explained by each component. These vectors are optimized jointly for each $k$ and are not ordered by the algorithm itself. D) The Grassmann distance between subspaces spanned by different linear measurements as a function of $k$ (Hamm & Lee, 2008). We show the distance between subspaces spanned by PCs obtained on two random halves of the training data (blue), between subspaces spanned by PCs and OLMs (red), and between subspaces spanned by PCs and a set of random measurements (green). See Appendix E for the definition of Grassmann distance. E) The (absolute) skewness value of the measurement on each component as a function of number of measurements ($k$, sorted post hoc based on the skewness value) for the PCs (blue) and OLMs (red). F) Distribution of mutual information between all pairs of measurement directions for PCA (blue), ICA (green), and OLM (red). G) Performance (on MNIST) and visualization of OLMs (olive, $\text{OLM}_{\text{seq}}$) obtained via a sequential version of the algorithm Section E.6. H) Performance as a function of measurement noise (on MNIST, $k=32$) and visualization of OLMs obtained with Gaussian noise ($\sigma=0.4$) added to the linear measurements during optimization (olive, $\text{OLM}_{\text{noise}}$), along with comparison to the regular OLMs (red) and PCs (orange).
@@
In this section, we present some qualitative and quantitative analyses to help interpret the difference between the linear measurement subspaces defined by PCs and OLMs. For the MNIST dataset, as expected, the first few PCs appear digit-like, and the measurement vectors contain increasing high-spatial frequency content as $k$ increases (Figure 4A, left). The OLMs, on the other hand, do not follow this coarse-to-fine (low to high frequency) ordering and are visually much more similar to each other (Figure 4A, right). This is reflected, quantitatively, in the measurement variance as a function of $k$, which falls rapidly for the PCs, but slowly for the OLMs (Figure 4C, left).
@@
Similar phenomena are observed for the CelebA dataset (Figure 4B, C). The PCs, known as “eigenfaces” (Sirovich & Kirby, 1987; Turk & Pentland, 1991), contain features that are geometrically aligned with features of the face, with increasing spatial frequency content. The variance of the measurements across PCs falls approximately exponentially. Although the OLMs also have a face-like appearance, they differ from the PCs. Each OLM vector has a detailed representation of fine facial features (eyes, nose, lips), and a coarse representation of the surrounding content (forehead, chin, hair). As in the MNIST case, all OLMs have similar spatial frequency composition. Consistent with this, the variance of the measurements falls slowly across the OLMs (Figure 4C, right).
@@
We can also characterize the linear measurements by analyzing the distribution of measurement values they produce in response to natural images (i.e., the measurement distribution). We found that OLMs differ from traditional methods in two key aspects: First, their measurement distributions are highly asymmetrical, as indicated by the significant non-zero skewness values (see Figure 4E). Second, unlike PCA or ICA, OLMs produce non-factorized measurement distributions: the measurements across axes are not fully independent. This dependency is reflected in the increased pairwise mutual information between different measurement directions observed for OLMs compared to those obtained with PCA or ICA (Figure 4F). This observation aligns with the intuition from the 2D example in Figure 1, where optimal measurements for non-Gaussian priors are not necessarily aligned with the axes.
Lastly, we also consider two modifications to the optimization procedure that result in different OLMs. First, in Figure 4G we present a set of OLMs that are sequentially organized – similar to PCs – using a greedy version of the algorithm (see Section E.6), where the measurement subspace for smaller $k$ is always contained within that of larger $k$. We found that while enforcing this sequential constraint results in a slight decrease in performance, the resulting OLMs exhibit more ordered spatial frequency content. Nonetheless, they remain distinct from the PCs both visually and when comparing subspace distances (not shown).
Second, in Figure 4H we examine the robustness of different measurement matrices to noise. For both PCs and OLMs, PSNR decreases as an increasing amount of Gaussian noise is added to the measurements. The OLMs can be made more robust to noise by injecting measurement noise during the OLM optimization (in this case, with a fixed standard deviation of $\sigma=0.4$). This results in a more noise-robust set of linear measurements ($\text{OLM}_{\text{noise}}$), accompanied by a modest performance loss in low-noise regimes. The noise-optimized OLMs are less smooth than those optimized without noise (Figure 4H). A more thorough investigation of noise robustness is worthwhile, but will require the development of a reconstruction algorithm that operates on noisy measurements.
@@ 3.5 Optimized linear measurements for perceptual loss
Figure 5: Optimal measurements for perceptual loss, as measured by the Structural Similarity Index Measure (SSIM, Wang et al. (2004)). A) Left: SSIM as a function of number of measurement $k$, for PC projection, PC, and OLM optimized for MSE and SSIM, all three using the denoiser prior. The number indicates the increase in SSIM from OLM optimized for MSE to OLM optimized for SSIM. Right: Scatter plot of the SSIM value for individual images, for OLM optimized for SSIM (x-axis) and MSE (y-axis), respectively. B) Example linear inverse solutions obtained from different measurement matrices (at $k=32$) combined with the denoiser prior. The numbers indicate the SSIM value for each individual image. C) Example measurement vectors from the OLM optimized for SSIM.
The previous sections described measurements optimized for MSE. However, our method is general, and can be used to optimize any differentiable objective function defined on the linear inverse estimates.
As an example, we computed a set of OLMs with respect to structural similarity index (SSIM) (Wang et al., 2004), a widely used perceptual image quality metric. Figure 5 shows results obtained using this set of optimal linear measurements. We obtained a noticeable improvement in mean SSIM for a range of $k$, and this improvement applies to the majority of the images in the test set. In Figure 5B, four example images are shown to highlight the visual improvement obtained by optimizing for the SSIM objective. However, it is worth noting that optimizing for SSIM does result in a lower average PSNR compared to the PCs, suggesting a trade-off between optimizing for perceptual quality versus MSE reconstruction, at least with our methods.
Lastly, in Figure 5C, the SSIM-optimized measurement vectors are visualized, which appear quite distinct from both the PCs and their MSE-optimized counterpart shown in Figure 4B. We further confirmed this difference by analyzing the measurement subspace, and found that the subspace distance between OLMs optimized toward different objectives is comparable to their distance from the PCs.
@@ 4 Related Work
Image Priors. Traditionally, prior models have been developed by combining constraints imposed by structural properties, such as translation or dilation invariance, with simple parametric forms, such as Gaussian, mixtures of Gaussian, or local Markov random fields. These models have been used for solving inverse problems with steady and gradual improvement in performance (Donoho, 1995; Simoncelli & Adelson, 1996; Moulin & Liu, 1999; Hyvärinen, 1999; Romberg et al., 2001; Şendur & Selesnick, 2002; Portilla et al., 2003a; Lyu & Simoncelli, 2008; Zoran & Weiss, 2012). More recently, methods such as VAEs (Kingma & Welling, 2013) and GANs (Goodfellow et al., 2020) have been developed to learn more complex image priors by taking advantage of the increased expressivity of deep neural networks. In the last few years, score-based diffusion models have emerged as the state-of-the-art method for learning sophisticated image priors (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Nichol & Dhariwal, 2021), as evidenced by the high quality of generated images obtained as draws from these priors. Diffusion models rely on the explicit relationship between the MMSE denoising solution and the score of the noisy image density (Miyasawa, 1961; Raphan & Simoncelli, 2011; Vincent, 2011). In addition to enabling unconditional sampling from the implicit prior, the power of that prior can be utilized to obtain high-quality solutions to inverse problems (Kadkhodaie & Simoncelli, 2020; Kawar et al., 2022; Wang et al., 2022; Chung et al., 2022; Zhang et al., 2024; Li et al., 2024). A closely related line of work known as plug-and-play (P&P) (Venkatakrishnan et al., 2013) used denoisers as regularizers for solving inverse problems. A number of recent extensions have used this concept to develop MAP solutions for inverse problems (Chan et al., 2016; Romano et al., 2017a; Zhang et al., 2017; Kamilov et al., 2017; Meinhardt et al., 2017; Chan et al., 2017; Mataev et al., 2019; Teodoro et al., 2019; Sun et al., 2019; Reehorst & Schniter, 2019; Pang et al., 2020; Sun et al., 2023).
@@
Optimized Linear Measurements. Other work has considered the problem of optimizing linear measurements directly. Weiss et al. (2007) observed that neither PCA nor random measurement fully takes into account the statistics of natural images. They attempted to optimize linear measurement using information-based criteria, but were not able to find measurements that outperformed PCA. Wu et al. (2019b) also proposed a method to find better measurement functions (i.e., both linear and nonlinear) in a compressed sensing framework by using the restricted isometry property (Candes, 2008) as an objective. They showed that the optimized measurement functions (both linear and nonlinear) are superior to simple random measurements, but did not compare their results with PCA. They also proposed an optimization method for finding additional structure in signals beyond sparsity (Wu et al., 2019a). Pineda et al. (2020) proposed a method to improve sampling efficiency in an fMRI setting through reinforcement learning. In a different line of work, Burge & Geisler (2015); Herrera-Esposito & Burge (2025) developed a method for finding linear measurements that are optimal for specific downstream tasks, such as estimating the motion presented in a video sequence.
@@ 5 Conclusion
We present a method for finding optimal linear measurements, using a nonlinear reconstruction method based on a learned prior embedded in a denoiser. We show numerically that the set of linear measurements found through our method results in better image reconstruction than PC measurement, despite the fact that the latter captures more signal variance. The key components of our method are: (i) a denoising diffusion model, which allows us to learn a complex prior underlying the datasets; (ii) a linear inverse algorithm that uses the diffusion prior to generate image reconstructions from linear measurements; and (iii) an end-to-end optimization procedure to find the measurement matrix that minimizes a loss function on the reconstructed images. We show that the optimal measurements are sensitive to both the statistics of the image dataset, and to the objective function for which they are optimized. Our results highlight the importance of accurately modeling the statistics of the signals to design efficient linear measurements.
One major limitation of our current work is that optimizing linear measurements through the iterative diffusion process is substantially more computationally expensive than computing PCs or generating random CS measurements. Additionally, the OLMs are obtained separately for each $k$, and are therefore not sequentially ordered like the principal components. We have restricted ourselves to linear measurements to maintain a direct connection to classical signal processing literature, as they are easier to understand and visualize. Despite this simplifying assumption, it is difficult to ascertain the relationship of the learned measurements to the diffusion prior, and thus to the reconstructed images. It is of interest to expand the current results to linear measurements with realistic noise models, or additional measurement constraints such as locality or non-negativity. Finally, developing reconstruction methods that couple DNN priors with more general optimized nonlinear measurements presents a challenging but enticing goal for future research.
@@ Appendix B Principal Components Analysis
To minimize $l(M)$, the $m_{i}$’s should be the eigenvectors associated with the largest eigenvalues of $XX^{T}$, which is the empirical covariance matrix. Note that this solution does not assume a Gaussian data distribution: For any data, the optimal (MMSE) linear projection is the projection onto the leading principal components of the covariance matrix.
When the data is in fact Gaussian distributed, linear projection onto the first $k$ PCs does achieve the minimal error possible, as shown in the bivariate example in Fig. 1A. For non-Gaussian distributions, nonlinear estimation can potentially improve performance.
@@ Appendix D Constrained Sampling Algorithm
The impact of these parameters on reconstruction performance, along with the number of conditional samples used to obtain the MMSE, is illustrated in the figure below.
Supplementary Figure 1: The effect of different parameters (step size $h$, injected noise $\beta$, stop criterion $\sigma_{\text{end}}$, and number of conditional samples used for averaging) of the algorithm on reconstruction performance, demonstrated using the OLM at $k=32$ on the MNIST dataset.
In general, the step size $h$ has a very small effect (note the scale of the y-axis). Achieving good reconstruction performance requires a larger amount of injected noise and a sufficiently small noise stop criterion. Additionally, averaging multiple conditional samples significantly improves performance, but the effect quickly diminishes as $n$ increases.